# UAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers

### v1

@@ Appendix A Mathematical Details
In this section, we provide the formal mathematical foundations underlying the uncertainty modeling and diagnostic analyses used in UAT-Lite. We first introduce a layer-wise variance decomposition that attributes predictive uncertainty across transformer depth under stochastic inference. We then provide a proof of the decomposition, clarify its assumptions, and relate Monte Carlo dropout to approximate Bayesian inference. Finally, we discuss practical and interpretive implications of layer-wise uncertainty attribution for calibration, intervention, and uncertainty-aware inference. These analyses are intended as diagnostic and explanatory tools that complement the empirical results in the main paper, rather than as claims of exact posterior inference.
@@ A.1 Layer-Wise Uncertainty Decomposition
To better understand how predictive uncertainty emerges within the transformer, we analyze its distribution across layers under stochastic inference.
@@
Table 5 presents representative layer-wise uncertainty decompositions for answerable and unanswerable SQuAD examples, illustrating how predictive uncertainty is distributed across transformer depth rather than manifesting solely at the output layer. We emphasize that these examples are illustrative diagnostics rather than statistical summaries, and are intended to provide qualitative insight into uncertainty propagation rather than population-level guarantees.
@@
In contrast, the answerable example exhibits a similar normalized uncertainty distribution but with a different internal structure. Mid-layer attention contributes more strongly to evidence aggregation (8.7%), while late-layer uncertainty reflects confident task-level reasoning rather than unresolved ambiguity.
@@ A.2 Proof of Theorem 1
We derive the layer-wise variance decomposition by applying the law of total variance recursively across transformer layers.
@@
Let denote the model’s final prediction (e.g., class probability or logit), and let denote the hidden representation at layer corresponding to the [CLS] token. Stochasticity is introduced via Monte Carlo dropout at inference time.
@@
where denotes dropout-induced noise at layer . We assume finite second moments for all noise distributions but do not assume Gaussianity or independence across layers.
Layer-wise variance contribution.
We define the variance contribution of layer as
Recursive variance decomposition.
Applying the law of total variance at the final layer yields
Applying the same decomposition recursively yields an additive expansion over layers:
@@
This decomposition provides an attribution of predictive uncertainty across layers under conditional expectations. While inter-layer dependencies may introduce higher-order interaction effects, the decomposition remains useful as an interpretable diagnostic of where uncertainty is amplified during inference. ∎
@@ A.3 Bayesian Interpretation of MC Dropout
The variability across predictions reflects epistemic uncertainty arising from model and data limitations, which is reducible with additional information.
@@ A.4 Implications of Layer-Wise Uncertainty
Decomposing predictive variance across transformer depth enables identification of where uncertainty is amplified during inference. As shown in Table 5, unanswerable SQuAD examples concentrate substantial uncertainty in late layers (9–11), consistent with ambiguity emerging during final reasoning stages. Answerable examples exhibit comparable late-layer contributions, but with uncertainty reflecting confident task-level reasoning rather than unresolved ambiguity.
@@
Layer-wise attribution suggests that calibration strategies could in principle target high-variance components. Our ablation study (Fig.˜4) shows that uncertainty introduced at the attention level yields a 21.1% improvement in calibration, while embedding-level uncertainty contributes negatively (-3.1%), supporting selective intervention rather than uniform stochasticity.
@@ A.5 Practical Inference-Time Procedure
Uncertainty estimates influence inference by modulating attention behavior under representation ambiguity and enabling confidence-aware decision shaping at the output level. All uncertainty mechanisms operate strictly at inference time and do not modify training dynamics or learned parameters.
@@ Appendix C Baselines and Metrics
We compare against BERT-base (110M parameters), MC Dropout (, ), Temperature Scaling (TS), and Deep Ensembles (5 independently fine-tuned models). Temperature scaling is treated as a strong post-hoc calibration baseline that operates purely at the output level and does not modify internal model computation.
Because TS rescales logits without affecting representations, it is compatible with uncertainty-aware inference and could in principle be applied on top of UAT-Lite. In this work, however, we report TS as a standalone baseline and do not include a stacked (UAT-Lite + TS) variant in the main tables, in order to isolate the effects of uncertainty-aware internal computation from output-level rescaling.
All baselines use the same pretrained backbone, fine-tuning recipe, and evaluation code unless otherwise stated.
@@
where predictions are partitioned into fixed-width confidence bins over . For each bin, denotes empirical accuracy and denotes the mean predicted confidence. Here denotes the number of confidence bins used for calibration evaluation, distinct from the number of Monte Carlo samples used for uncertainty estimation. Lower ECE indicates better calibration.
@@
We evaluate SQuAD 2.0 on the answerability prediction task, where the model must determine whether a question is answerable given the context. This binary classification task is well-suited for calibration analysis. We report answerability prediction accuracy alongside ECE in Table 1 . Note that span-level extraction (predicting answer start/end positions) is a separate task and orthogonal to our uncertainty calibration contribution. Our focus is on calibrated confidence scores for the answerability decision, not optimizing span extraction performance.
@@
where smaller (or negative) values indicate improved robustness under shift. Second, we report a calibration robustness score defined as the average ECE across ID and OOD settings:
@@
We additionally report selective coverage at fixed confidence thresholds and the Area Under the Risk–Coverage curve (AURC; lower is better). Confidence thresholds are selected on the validation set and held fixed at test time for all methods.
@@
The MC Dropout baseline uses a uniform dropout rate of across all layers, consistent with prior calibration studies. While more aggressive or component-specific dropout tuning could further improve MC Dropout performance, such tuning increases implementation complexity and reduces comparability with standard baselines. In contrast, UAT-Lite explicitly leverages structured uncertainty within the attention mechanism rather than relying solely on increased stochasticity. This comparison isolates the effect of uncertainty-aware attention rather than raw dropout intensity.
@@
To explicitly disentangle the effect of structured stochasticity from uncertainty-aware attention, we additionally evaluate an MC Dropout baseline using the same component-specific dropout rates as UAT-Lite (0.1 embeddings / 0.2 attention / 0.3 feed-forward), but without attention modulation. As shown in Table 1, this control does not improve calibration relative to uniform MC Dropout or the BERT baseline, indicating that the gains of UAT-Lite arise from uncertainty-weighted attention rather than tuned dropout alone.
@@
Within each table, all methods are evaluated using the same data splits, evaluation code, confidence definitions, and Monte Carlo budget (where applicable). ECE is computed with fixed-width bins for all methods. Temperature scaling (and UAT-Lite +TS) fits a single scalar temperature on the validation set only, and is applied at test time without any access to test labels. OOD calibration metrics (e.g., ECE and Robustness) are computed using the identical pipeline on ID and OOD sets to avoid binning or confidence-definition mismatch.
@@
Main-paper results (general calibration, distribution-shift robustness, and selective prediction) are reported as mean standard deviation over five random seeds. Appendix diagnostic analyses—including qualitative attention visualizations, representative layer-wise uncertainty decompositions, and linguistic or adversarial case studies—are reported using a single seed to reduce computational cost and are explicitly labeled as single-seed in the corresponding captions. Deterministic baselines exhibit negligible variance because inference is fully deterministic when evaluated from a fixed fine-tuned checkpoint reused across seeds.

### v2

@@ Appendix A Mathematical Details
In this section, we present the formal mathematical foundations underlying the uncertainty modeling and diagnostic analyses used in UAT-Lite. We first introduce a layer-wise variance decomposition that attributes predictive uncertainty across transformer depth under stochastic inference. We then provide a proof of the decomposition, clarify its assumptions, and relate Monte Carlo dropout to approximate Bayesian inference. Finally, we discuss the practical and interpretive implications of layer-wise uncertainty attribution for calibration, intervention, and uncertainty-aware inference. These analyses serve as diagnostic and explanatory tools that complement the empirical results in the main paper, rather than as claims of exact posterior inference.
@@ A.1 Layer-Wise Uncertainty Decomposition
To better understand how predictive uncertainty arises within the transformer, we analyze its distribution across layers under stochastic inference.
@@
Table 3 presents representative layer-wise uncertainty decompositions for answerable and unanswerable SQuAD examples, illustrating how predictive uncertainty is distributed across transformer depth rather than appearing only at the output layer. We emphasize that these examples are illustrative diagnostics rather than statistical summaries and are intended to provide qualitative insight into uncertainty propagation rather than population-level guarantees.
@@
In contrast, the answerable example exhibits a similar normalized uncertainty distribution but a different internal structure. Mid-layer attention contributes more strongly to evidence aggregation (8.7%), while late-layer uncertainty reflects confident task-level reasoning rather than unresolved ambiguity.
@@ A.2 Derivation of Proposition 1
We derive a diagnostic layer-wise variance decomposition by recursively applying the law of total variance across transformer depth. The purpose of this derivation is interpretive: it provides a layer-indexed summary of where predictive variance arises or amplifies during stochastic inference.
@@
Let denote the model’s final prediction (e.g., a class probability or logit), and let denote the hidden representation at layer corresponding to the [CLS] token. Stochasticity is introduced through Monte Carlo dropout at inference time.
@@
where denotes dropout-induced noise at layer . We assume finite second moments so that all conditional variances below are well defined.
Recursive application of total variance.
Applying the law of total variance at the final layer gives
The same identity can then be applied recursively to the second term by conditioning on progressively earlier hidden representations . This yields a recursive decomposition of predictive variance across depth.
Diagnostic layer-wise contributions.
Motivated by this recursion, we define the layer-wise diagnostic contribution of layer as
These quantities summarize the predictive variability associated with the stochasticity introduced at each layer, conditioned on the representation propagated from preceding layers.
@@
Under this construction, predictive variance admits a recursive layer-indexed decomposition, and the normalized values
provide an interpretable summary of how uncertainty is distributed across depth. We emphasize that these contributions are intended as diagnostic attributions: they are useful for identifying where uncertainty concentrates or amplifies during inference, but they should not be interpreted as unique causal effects or as a novel variance identity. ∎
@@ A.3 Bayesian Interpretation of MC Dropout
Variability across predictions reflects epistemic uncertainty arising from model and data limitations and is, in principle, reducible with additional information.
@@ A.4 Implications of Layer-Wise Uncertainty
Decomposing predictive variance across transformer depth enables identification of where uncertainty is amplified during inference. As shown in Table 3, unanswerable SQuAD examples concentrate substantial uncertainty in late layers (9–11), consistent with ambiguity emerging during final reasoning stages. Answerable examples exhibit comparable late-layer contributions, but the uncertainty reflects confident task-level reasoning rather than unresolved ambiguity.
@@
Layer-wise attribution suggests that calibration strategies could, in principle, target high-variance components. Our ablation study (Fig.˜5) shows that uncertainty introduced at the attention level yields a 21.1% improvement in calibration, while embedding-level uncertainty contributes negatively (-3.1%), supporting selective intervention rather than uniform stochasticity.
@@
Scope and limitation (per-layer recomputation).
While we analyze where uncertainty concentrates across depth via layer-wise variance attribution (Table 3), our UAT-LITE intervention computes token uncertainty once (from stochastic forward passes) and reuses it across layers for efficiency. Recomputing token uncertainty after each layer would require defining uncertainty over mixed representations after contextualization and would increase inference cost roughly in proportion to the number of layers; we leave such layer-wise recomputation to future work.
@@ A.5 Practical Inference-Time Procedure
Uncertainty estimates influence inference by modulating attention under representation ambiguity and enabling confidence-aware decision shaping at the output level. All uncertainty mechanisms operate strictly at inference time and do not modify training dynamics or learned parameters.
@@ Appendix C Baselines and Metrics
We compare against BERT-base (110M parameters), MC Dropout (, ), TS, and Deep Ensembles (5 independently fine-tuned models). TS is treated as a strong post-hoc calibration baseline that operates purely at the output level and does not modify internal model computation. Because TS rescales logits without affecting representations, it is compatible with uncertainty-aware inference and can be applied on top of UAT-Lite. Accordingly, we report both TS as a standalone baseline and the stacked variant UAT-Lite +TS to assess composability. For UAT-Lite +TS, the temperature is fit on the development set using the MC-mean logits produced by UAT-Lite and applied at test time. To isolate the effect of uncertainty-aware internal computation, we compare UAT-Lite against non-stacked baselines (e.g., Base, MC Dropout, and TS), and treat UAT-Lite +TS as an explicit compositional setting rather than the core mechanism itself. All baselines use the same pretrained backbone, fine-tuning recipe, and evaluation code unless otherwise stated.
@@
where predictions are partitioned into fixed-width confidence bins over . For each bin, denotes empirical accuracy, and denotes mean predicted confidence. Here, denotes the number of confidence bins used for calibration evaluation, which is distinct from the number of Monte Carlo samples used for uncertainty estimation. Lower ECE indicates better calibration.
@@
We evaluate SQuAD 2.0 on the answerability prediction task, in which the model determines whether a question is answerable given the context. This binary classification setting is well suited to calibration analysis. We report answerability prediction accuracy alongside ECE in Table 1. Span-level extraction (predicting answer start and end positions) is a separate task and is orthogonal to our uncertainty calibration contribution. Our focus is on calibrated confidence scores for the answerability decision rather than on optimizing span extraction performance.
@@
where smaller (or negative) values indicate greater robustness under shift. Second, we report a calibration robustness score defined as the average ECE across ID and OOD settings:
@@
We additionally report selective coverage at fixed confidence thresholds and the Area Under the Risk–Coverage Curve (AURC; lower is better). Confidence thresholds are selected on the validation set for each method and then held fixed at test time. We also report admitted counts () at fixed thresholds to make score-rescaling effects (e.g., under TS) explicit.
@@
The MC Dropout baseline uses a uniform dropout rate of across all layers, consistent with prior calibration studies. Although more aggressive or component-specific dropout tuning could further improve MC Dropout performance, such tuning increases implementation complexity and reduces comparability with standard baselines. In contrast, UAT-Lite explicitly leverages structured uncertainty within the attention mechanism rather than relying solely on increased stochasticity. This comparison isolates the effect of uncertainty-aware attention rather than raw dropout intensity.
@@
To disentangle the effect of structured stochasticity from uncertainty-aware attention, we additionally evaluate an MC Dropout baseline using the same component-specific dropout rates as UAT-Lite (0.1 embeddings / 0.2 attention / 0.3 feed-forward), but without attention modulation. As shown in Table 1, this control does not improve calibration relative to uniform MC Dropout or the BERT baseline, indicating that the gains of UAT-Lite arise from uncertainty-weighted attention rather than tuned dropout alone.
@@
Within each table, all methods are evaluated using the same data splits, evaluation code, confidence definitions, and Monte Carlo budget (where applicable). ECE is computed with fixed-width bins for all methods. Temperature scaling (and UAT-Lite +TS) fits a single scalar temperature on the validation set only and is applied at test time without access to test labels. OOD calibration metrics (e.g., ECE and Robustness) are computed using the same pipeline on ID and OOD sets to avoid binning or confidence-definition mismatch.
@@
Main-paper results (general calibration, distribution-shift robustness, and selective prediction) are reported as mean standard deviation over five random seeds. Appendix diagnostic analyses, including qualitative attention visualizations, representative layer-wise uncertainty decompositions, and linguistic or adversarial case studies, are reported using a single seed to reduce computational cost and are explicitly labeled as single-seed in the corresponding captions. Deterministic baselines exhibit negligible variance because inference is fully deterministic when evaluated from a fixed fine-tuned checkpoint reused across seeds.